OpenStudy (goformit100):

The length x of a rectangle is decreasing at the rate of 3 cm/minute and the width y is increasing at the rate of 2cm/minute. When x =10cm and y = 6cm, find the rates of change of (a) the perimeter and (b) the area of the rectangle.

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

hmm... tricky... \[\huge P = 2l + 2w\]\[\huge A = lw\]

terenzreignz (terenzreignz):

\[\Large \frac{dl}{dt}=-3\]\[\Large \frac{dw}{dt}=2\]

OpenStudy (goformit100):

Thank you Sir

terenzreignz (terenzreignz):

But my instincts say... \[\huge \frac{dP}{dt}=\frac{\partial P}{\partial l}\cdot{\frac{dl}{dt}}+\frac{\partial P}{\partial w}\cdot\frac{dw}{dt}\] For area, just replace P with A, accordingly.

terenzreignz (terenzreignz):

oh sorry, I guess replace l with x and w with y.